OpenStudy (anonymous):

What are the factors of 3x2 + 10x + 3? a.) (3x − 1)(x − 3) b.) (3x − 3)(x − 1) c.) (3x + 1)(x + 3) d.) (3x + 3)(x + 1)

OpenStudy (anonymous):

\[3x^2 + 10x + 3 = 3x^2 + 9x +x + 3 = 3x(x+3)+1(x+3) = (x+3)(3x+1)\]Hence (x+3)(3x+1) are the required factors of the given expression.

OpenStudy (anonymous):

Thank you lots.

OpenStudy (anonymous):

Hence option c.) (3x + 1)(x + 3) is correct

OpenStudy (anonymous):

Yes